OpenStudy (lgbasallote):

could someone show me how to solve multiple integral? im gonna write one of the examples given to us before and i would appreciate if someone could show me the step by step to solve it ^_^

OpenStudy (lgbasallote):

\[\LARGE \int\limits_{0}^{1}\int\limits_{0}^{z}\int\limits_{0}^{y} (z^{2}xy) dx dy dz\] i think it's a simple one..just wanna see the process

OpenStudy (experimentx):

first integrate x  put that 1/2 in place of x, and remove one integration sign and dx